this copy of the exam at home. You are free to use your textbook and notes while completing the exam. You MUST turn in the co number 1 thank you.
Answer #1

At each half-life, the number of atoms is divided by 2, then:

3 half-lives = 20000/2/2/2 = 2500 atoms

6 half-lives = 20000/2/2/2/2/2/2 = 312.5 atoms

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.